OpenStudy (anonymous):

What is the probability that a randomly thrown dart that lands within the rectangle lands within a shaded region? All of the circles are congruent, and the diameter of each circle is 28 cm. A. 0.112 B. 0.196 C. 0.697 D. 0.785

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is the ratio of the area of the circle to the area of the rectangle

OpenStudy (misty1212):

a picture would be really really helpful

OpenStudy (anonymous):

ugh im trying to draw it buts its hard

OpenStudy (anonymous):

there we go loll

OpenStudy (anonymous):

??

OpenStudy (misty1212):

ok now we can do it

OpenStudy (anonymous):

what are the side lenghts

OpenStudy (misty1212):

you are actually told the diameter of each circle? it is 28 right (not that it matters one bit, just checking)

OpenStudy (anonymous):

yess

OpenStudy (misty1212):

the base of your rectangle is therefore \(28\) and the height is \(4\times 28=112\) so the area of the rectangle is ?

OpenStudy (misty1212):

i get \[28\times 112=3136\] as the area

OpenStudy (anonymous):

then you would subtract the area of the four circles from the area of the rectangle

OpenStudy (misty1212):

each circle has radius \(14\) so area \(4*14^2\times \pi\) or about \(2463\)

OpenStudy (anonymous):

theres no side lengths though thts the problem

OpenStudy (misty1212):

lol no dear but we know it anyway, because the diameter is 28

OpenStudy (misty1212):



OpenStudy (anonymous):

and i get 673?

OpenStudy (misty1212):

hmm no i get \[28\times 112=3136\]

OpenStudy (misty1212):

for the total area that is

OpenStudy (anonymous):

no 3136-2463

OpenStudy (misty1212):

ooh ok i see yes

OpenStudy (misty1212):

so \[\frac{673}{3136}\] for your answer

OpenStudy (anonymous):

thats not an answer

OpenStudy (misty1212):

lol that is because i read it wrong the second time it is not "outside of the circles" but "inside of the circles" just \[\frac{2463}{3136}\]

OpenStudy (misty1212):

hmm also not quite lets do the numbers again \[\frac{4\times 14^2\times \pi}{28\times 112}\]

OpenStudy (misty1212):

also known as \(\frac{\pi}{4}\)

OpenStudy (anonymous):

thanks!!